OpenStudy (anonymous):

Solve cos X tan X - sin ^2 X=0 for 0 to 360 degrees...??? A. 0,pi/4,pi B. 0,pi/2 C. 0,pi/2,pi D. 0,pi, 3pi/2 I think the answer is D

OpenStudy (danjs):

ok

OpenStudy (danjs):

\[\cos(x)\frac{ \sin(x) }{ \cos(x) } - \sin^2(x) = 0\]

OpenStudy (danjs):

so \[\sin(x) - \sin^2(x) = 0\]

OpenStudy (danjs):

pull out a sin(x) \[\sin(x)* [ 1-\sin(x)] = 0\]

OpenStudy (danjs):

so you have 2 equations sin(x) = 0 and sin(x) = 1

OpenStudy (danjs):

Recall, the sin value on the unit circle is the y coordinate of any point on the unit circle.

OpenStudy (danjs):

When on the circle, is the Y value 0 or 1?

OpenStudy (danjs):

When the angle is 0 degrees, the point on the unit circle is (1,0)

OpenStudy (danjs):

so the sin(x) = 0, when the angle is 0 degrees

OpenStudy (danjs):

Likewise, if the angle is 180 degrees, you are at the point (-1,0), the y value here is also zero, so sin(180) also is zero,

OpenStudy (danjs):

Now for sin(x) = 1 The Y value is 1, when you are at 90 degrees, you are at point (0,1) on the unit circle

OpenStudy (misty1212):

how is it from 0 to 360 if your answers have \(\pi\) in them ? shouldn't it be \(0\) to \(2\pi\)?

OpenStudy (anonymous):

yes

OpenStudy (danjs):

ok, 0 degrees is 0 radians 90 degrees is pi/2 radians 180 degrees is pi radians

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

and I need help with one more question

OpenStudy (danjs):

For Reference...

OpenStudy (anonymous):

so the right answer is c

OpenStudy (danjs):

yes, because at 3/2 pi or 270 degrees, the y value of the point is -1 on the y axis y = sin(x) = -1 which is not +1

OpenStudy (danjs):

Sin(x) has to be either 0, or +1 Which occurs when the y value on that circle is 0 or +1 Which occurs at 0, pi/2, and pi angles

OpenStudy (anonymous):

ok I got it and you explained me very well

OpenStudy (danjs):

cool

OpenStudy (anonymous):

Find the ordered triple that represents the vector from A(-3,5,6) to B(-6,8,6). Then find the magnitude of AB? A. (3,-3,0); 4.24 B. (-9,13,12); 19.85 C. (-3,3,0); 4.24 D. ( -9,3,0); 9.49

OpenStudy (danjs):

this one is pretty easy...

OpenStudy (anonymous):

really how can you solve this problem

OpenStudy (danjs):

AB = <(Bx - Ax) , (By - Ay), (Bz - Az)>

OpenStudy (danjs):

The vector is that

OpenStudy (danjs):

The magnitude of vector AB = <x , y, z> is \[\left| AB \right| = \sqrt{x^2 + y^2 +z^2} = \sqrt{(Bx - Ax)^2 + (By - Ay)^2 + (Bz - Az)^2}\]

OpenStudy (danjs):

so AB = < (-6-(-3)) , (8-5) , (6-6)>

OpenStudy (danjs):

AB = <-3, 3, 0> \[\left| AB \right| = \sqrt{(-3)^2 + 3^2 + 0^2}\]

OpenStudy (anonymous):

So the right answer is c

OpenStudy (danjs):

um